OpenStudy (8bitfox):

I need help please I'm stuck, Use commutative property of multiplication to solve: 2.30b + 1.85(23-b) = 46.60 I'm so confused I'm learning about cost problems and I simplified the equation to there now they want me to distribute and now I'm confused

OpenStudy (studygurl14):

distribute the 1.85 to the stuff in parenthesis. For example: 2(2 + x) = 4 + 2x

OpenStudy (8bitfox):

so 1.85(23-b) = 42.55 +23b?

OpenStudy (studygurl14):

Almost. It would be 42.55 - 1.85b

OpenStudy (8bitfox):

Oh I see thank you

OpenStudy (studygurl14):

I guess I should've used different numbers for my example... 2(3 + x) = 6 + 2x

OpenStudy (8bitfox):

so now Im down to 2.30b + 42.55 - 23b (im guessing I have to add the like terms now)

OpenStudy (studygurl14):

2.3b + 42.55 - 1.85b = 46.6

OpenStudy (8bitfox):

oh okay I think i got it from here thank you <3

OpenStudy (8bitfox):

so in the end b = 8.1 yes?

OpenStudy (studygurl14):

no

OpenStudy (studygurl14):

2.3b - 1.85b = 46.6 - 42.55 0.45b = 4.05 4.05 / 0.45 = 9 b = 9

OpenStudy (8bitfox):

oh I see I made a mistake simplifying Thank you I can now close this question

OpenStudy (studygurl14):

:)